### v3

# Neural Attention Search

Abstract
We present Neural Attention Search (NAtS), an end-to-end learnable sparse transformer that automatically evaluates the importance of each token within a sequence and determines if the corresponding token can be dropped after several steps. To this end, we design a search space that contains three token types: (i) Global Tokens will be preserved and queried by all the following tokens; (ii) Local Tokens survive until the next global token appears; and (iii) Sliding Window Tokens have an impact on the inference of a fixed size of the next following tokens. Similar to the One-Shot Neural Architecture Search approach, this token-type information can be learned jointly with the architecture weights via a learnable attention mask. Experiments on both training a new transformer from scratch and fine-tuning existing large language models show that NAtS can efficiently reduce the KV cache size and the inference costs for the models while maintaining the models’ performance.
1 Introduction
The ability to understand and infer from long-context information is crucial for many tasks such as long document summarization [91] and question answering [19, 46], code generation [33, 54] or multi-round dialogues [88]. Thanks to the ability to query the information from any position of the historical sequence, transformer-based large language models [10, 13, 30, 32, 40] extend their context length up to millions of tokens.
However, querying information from historical sequences requires a complexity of $\mathcal{O}(\mathit{L}^{2})$ w.r.t. the input sequence length $\mathit{L}$. KV caching could reduce this time complexity to $\mathcal{O}(\mathit{L})$ by storing all the historical KV values. Nevertheless, with the increasing model size of recent LLMs, even the $\mathcal{O}(\mathit{L})$ time-wise and memory-wise complexity could become a bottleneck during inference time.
Indeed, not all the tokens in a sequence are equally important [44]. Many of them are redundant and do not contribute to the final output. Humans can recognize this information without pre-defined fixed rules and summarize or discard the context information into much smaller content. Transformers could also learn this ability implicitly: Many tokens in the attention map might only have very low weights [93] and only have little influence on the final predictions. However, as the transformer learns this information implicitly, we might not know how the important tokens would be distributed in the context. Selecting these tokens and recognizing the attention distributions might require extra human experts’ knowledge by either looking at the attention maps [25, 56, 92, 93] or applying specific fixed rules [12, 15, 16, 29, 82]. Since this knowledge is already contained in the transformer models, we could also ask the model to evaluate the importance of each token and learn to predict the optimal type for the given input tokens automatically.
Unlike prior works that rely on human expertise or predefined rules to identify important tokens [14, 25, 26, 29, 52, 82, 83, 93], we propose a novel approach to evaluate the importance of each token by assigning different roles to each of the tokens. For example, some tokens will be preserved until the end, while other tokens might only survive for a short amount of time. These roles measure the importance of each token and determine if they would survive within the next few tokens. Rather than pre-define a set of fixed rules for each token, we ask the model to learn this information automatically. Finding the optimal role for each token is similar to the process of neural architecture search [21, 79], where an optimal architecture is searched by the optimizer for a given task. Thereby, in this work, we jointly optimize the choice of each token and the model weights by constructing a learnable attention mask. Our approach is implicitly motivated by the one-shot neural architecture search approach [20, 53, 66], where the model parameters and architecture parameters are jointly optimized during the search process. However, our approach, Neural Attention Search (NAtS), searches for the optimal token roles jointly with the attention weights.
Our contributions are as follows:
  1. 1.
    We propose Neural Attention Search (NAtS), an end-to-end learnable sparse attention framework that automatically searches for the optimal roles for each token in the input sequence.
  2. 2.
    We introduce different token roles in our search space that can be later combined to construct a learnable attention mask and then jointly optimized with the model weights in NAtS.
  3. 3.
    We show that NAtS  could efficiently reduce the KV cache required during inference time while maintaining most of the models’ performance.
By automatically learning to focus on the most relevant information, NAtS paves the way for more efficient and scalable inference with LLMs in long-context applications.
2 Background and Related Work
2.1 Attention maps for transformers
Transformers [77] computes the correlation between different tokens by mapping the input sequences into three variables, $Q$, $K$ and $V$. An attention map $A$ is first computed by pairing $Q$ and $K$ and then scaled and normalized with a softmax function. Finally, this attention map $A$ is multiplied by $V$. Additionally, a mask $M$ is attached to the Attention Map to guide the attention to focus only on certain tokens. This can be expressed as
$A=\frac{QK^{T}}{\sqrt{\mathit{d}_{head}}},\hskip 28.45274ptO=\mathrm{softmax}(A+M^{add})V$ (1)
The additive attention mask $M^{add}_{i,j}\in\{-\inf,0\}$ controls if the transformer needs to construct the correlation between $Q_{i}$ and $K_{j}$. A widely used mask $M^{add}$ used in transformers is a causal mask [77], where the upper parts of the mask are filled with $\inf$; each token can only query the token information before it.
Computing the attention map of a sequence with length $\mathit{L}$ requires a complexity of $\mathcal{O}(\mathit{L}^{2})$ for both time and memory. Techniques such as KV cache lower this complexity to the $\mathcal{O}(\mathit{L}_{\text{cache}})$ during inference time, with $\mathit{L}_{\text{cache}}$ being the length of the KV cache. However, as an expense, the cost of storing the KV cache gradually becomes unnegligible with the increasing size of large language models and the context length they could handle. Therefore, lots of work has been proposed to reduce the storage and computation overheads during the inference time [50].
To reduce the computational costs of an LLM on long context information, LLMLingua [64] trains another smaller model to compress the input texts. Many other studies work on reducing the KV cache size by deciding which tokens to evict or preserve with the past attention maps. Recent works such as H2O [93], Scissorhand [56], FastGen [29], KeyFormer [6], SnapKV [52], PyramidKV [14], AadaKV [25], ChunkKV [55], and CriticalKV [26] all apply different rules to identify the important of each token and remove all the remaining tokens. Additionally, instead of simply removing the unimportant tokens, we can also merge them into the existing tokens [78, 92].
Other works do not aim at reducing the KV cache size, but only select a subset of KV caches to compute the attention outputs, including Quest [74], TokenButler [8], AttentionPredictor [85], and XAttention [84]. However, these works still need to maintain the entire KV cache in the GPUs. This still results in a huge GPU memory usage overhead. To alleviate this issue, some other works like InfLLM [81], HIP [49], and ShodwKV [73] offload the KV cache to CPUs and only load the KV caches that are relevant for the current token predictions to GPUs.
While these existing methods offer various ways to reduce KV cache size, they often rely on inflexible predefined rules and potentially inaccurate heuristics based on past attention maps. NAtS, in contrast, introduces a novel approach that learns to dynamically assign token roles during inference, enabling a more adaptive and efficient use of the KV cache. By treating token role assignment as an optimization problem, NAtS leverages principles from neural architecture search to jointly optimize token roles and model weights, leading to a more flexible and powerful attention mechanism.
2.2 Sparse Attention
Sparse attention works on computing only a fraction of the attention map to reduce the computation costs and KV cache sizes with either pre-defined fixed patterns, including Sparse Transformers [16], Longformer [12], StreamingLLM [82], and DuoAttention [83]. Alternatively, this can be adjusted by special embedding tokens such as Longcoder [33] and SepLLM [15]. Additionally, one can also determine the types of sparse attention with a set of proxy values, including MInference [41], FlexPrefill [48], SeerAttention [28], SpargeAttention [90], and XAttention [84].
In contrast to the fine-tuning-free approaches above, we can also fine-tune the target model to achieve the desired sparsity. This includes Adaptively Sparse Attention [9] and Dynamic Memory Compression [62]. However, both approaches rely on expensive cumulative productions during training time, making their approach less applicable to long-context scenarios. LLMLingua2 [64] trains another model to compress the target texts. Additionally,  Kim et al. [44] asks the model to learn to drop the tokens with lower attention scores. Landmark Attention [61] inserts landmark tokens after certain time steps and applies these landmark tokens to summarize the information of the tokens before that landmark token and recover. SPARSEK [58] only selects a constant number of KV pairs for each query by introducing a differentiable SparseK operator. Mixture-of-Depth [70] only passes a limited number of tokens to each transformer layer. Native Sparse Attention [89] and MoBA [59] all select a subset of the KV values to compute the attention masks and therefore reduce the computational overhead. However, these approaches still require preserving the entire KV cache values and only reducing computational costs during decoding stages.
MoA [27] proposes to search for the optimal sparse attention for different heads, which share similar ideas to  NAtS. However, MoA’s search space only contains streaming LLM.This head-level search space of MoA is much smaller than our token-level search space, which might restrict the model’s expressibility under even smaller budgets.
2.3 Neural Architecture Search
Designing a neural architecture for a specific task might require a lot of trial and error. Neural Architecture Search (NAS) automates this process by searching in a pre-defined search space [22]. Previous work on NAS mainly focused on searching within a discrete search space by sampling a new architecture from the search space, evaluating its performance, and updating the optimizers [95, 96]. One-Shot NAS [66] approaches instead share the same weights of the operators with respect to all the architectures in the search space. This allows to jointly optimize the architectures and weights. DARTS [53] and GDAS [20] further relax the discrete search space into continuous values to optimize the architecture parameters with gradient descent. The One-Shot NAS approach allows the optimizers to efficiently search for the optimal architecture within a huge search space. Similarly, NAtS  has multiple options for each token as the search space and is able to search for the optimal token types jointly with the model weights. However, unlike One-Shot NAS approaches that consider the optimization problem as a bilevel optimization problem and optimize the model weights and architecture weights alternatively, we optimize the token state information and model weights within one forward and backward pass. This is similar to a mixture-of-experts (MOE) model [24, 71]. However, instead of distributing data across all experts uniformly, we only select one expert for each token and assign all data to that expert.
3 The NAtS Approach
In this section, we first introduce all the candidate token types in our search space. We then show that we can construct a learnable attention mask with the choice of each token type. Finally, we can efficiently reduce the KV cache size by dropping the unnecessary tokens during inference.
3.1 Search Space Design
Not all the tokens in a sequence are equally important. Just like a paragraph is composed of multiple sentences, a sequence can be divided into multiple sub-sequences containing different information; some tokens might only be required within these sub-sequences. Hence, we design a search space for each token’s role within the sequence and ask the model to automatically learn the optimal role for each token in the sequence.
We first define Global Tokens as tokens containing important information that need to be preserved for the following predictions. Liu et al. [56] and Zhang et al. [93] showed that only a small fraction of the tokens contribute to most of the attention scores for self-attention computation. These tokens need to be preserved for models to recall the global information that helps to predict the next token. In vanilla transformer models, all the tokens are Global Tokens.
Global Tokens will not be evicted during inference time. Therefore, we should maintain as few Global Tokens as possible to ensure inference efficiency. Each Global Token should not only preserve the information within that position. Ideally, it should also be able to summarize the information from previous sequences [15, 33]. Hence, we split the entire sequence with the Global Tokens into multiple sub-sequences, with each sub-sequence ending with one Global Token. Each Global Token only needs to summarize the information from its sub-sequences and the previous Global Tokens.
Local Tokens only survive until the next Global Token appears. Therefore, models will have the full attention within each sub-sequence to summarize the local sub-sequence information into the Global Token located at the end of the sub-sequence while being sparse within the input sequence.
Only the Global Tokens and Local Tokens might control the sparsity at a low granularity level. E.g., assuming that one input sequence is highly localized, each token only has a high correlation with itself or a few neighboring tokens. In this case, they are all similar and are assigned with the same token type. However, none of the Global Token and Local Token could sparsify this attention map efficiently: if all the tokens are classified as Local Tokens, the input sequence will only be considered as one single subsequence, and all the Local Tokens will be equivalent to the Global Tokens.
Hence, we introduce Sliding Window Token. Sliding Window Tokens will only be preserved for the next W time steps and were previously considered as one of the most popular sparse attention approaches [16, 29, 82, 93].
(a) A comparison between different casual attention maps. LABEL:fig:FullAtt: The full attention map, where each token is connected to the tokens before it. LABEL:fig:LocalAtt: the local attention with sliding windows 3, every token will only get access to the information of the 3 tokens ahead. LABEL:fig:LongFormer Longformer, besides the local attention, the first, 6th and 9th tokens are the pre-defined global tokens. LABEL:fig:LearnedAttention: NAtS dynamically learns the optimal role for each token and constructs a learnable mask based on the tokens’ roles.
In contrast to other causal attention maps, Figure LABEL:fig:LearnedAttention illustrates an exemplary attention mask constructed by the choices of different token types. In this case, we define the sliding window size as 4. Token 1, 4, 10 act as Global Tokens; Tokens 2, 6, 7, 8, 11 are Local Tokens; Token 3, 5, 9, 11 are Sliding Window Tokens. The Global Tokens splits the entire sequence into three subsequences that end at 4, 10, and the last index accordingly. Hence, Token 2 will only be required by Token 3 and 4. This rule applies the same for Token 6, 7, and 8, where they only interact with the tokens within the same subsequence. For the sliding window tokens, only the next three tokens could query their information, regardless of whether these tokens belong to the same sub-sequence. Only 6 out of 12 tokens are involved during inference time to make the next token prediction.
Combining different token types could already cover most of the sparse attention variables. For instance, assigning Global Tokens to the first few tokens and Sliding Window Tokens to all the other tokens results in streaming LLMs [82] and MoA [27]. If we further set all the tokens in some heads as Global Tokens, we will get DuoAttention [81]. Assigning different tokens as Global Tokens results in the Vertical patterns [41] that appear in many Token eviction strategies [52, 56]. Since different layers or heads might have different numbers of global tokens, patterns like pyramidkv [14] and adakv [25] can be easily derived from our search space. For a transformer network with $\mathit{N}$ layers and $\mathit{H}$ attention heads that needs to process one sequence of length $\mathit{L}$, given the three token types defined above, there are $3^{\mathit{L}\times\mathit{H}\times\mathit{N}}$ possible configurations in our search space. It is prohibitive to check all the combinations in such a huge search space, with the increasing sequence length that a model receives. Here, we propose an end-to-end approach to search for the optimal token types together with the network weights jointly.
3.2 Searching for the Optimal Token Types
Searching for the optimal token types within a sequence is similar to searching for the optimal architectures for a given task in neural architecture search [22, 79]. Following GDAS [20], we apply the Gumbel-Softmax trick [39, 60] to sample from the discrete search space. The Gumbel-Softmax trick allows the gradient to be backpropagated through the discrete choice of the token types.
Specifically, we first use a linear layer (which we call Attention Score Layer) that maps each the input tensor for an attention layer $X\in\mathbb{R}^{\mathit{d}}$ to the likelihood for each option: $\alpha\in\mathbb{R}^{(\mathit{H}*\mathit{N_{\text{opts}}})}=\text{Linear}(\mathbf{x})$, where $\mathit{H}$ is the number of KV heads and $\mathit{N_{\text{opts}}}$ is the number of options in the search space. The type $\alpha$ for each token is then sampled from a Gumble-Softmax function based on the likelihood values.
[…]
$O=\frac{e^{A}\odot M^{\text{mul}}}{\sum_{j}e^{A_{.,j}}\odot M^{\text{mul}}_{.,j}}V$ (2)
The attention mask columns for Global Tokens and Sliding Window Tokens can be directly constructed since they will survive for a fixed number of steps. However, the mask for Local Tokens $M^{L}_{i,j}$ is controlled by both the distribution from Local Tokens and Global Tokens as Local Tokens will survive until the next Global Token appears. In other words, to make $M^{L}_{i,j}(j>i)$ a valid value, no Global Token should appear between $i$ and $j$.
Formally, the attention masks can be created as follows:
$M^{G}_{i,j}$ $=1$ (3)
$M^{SW}_{i,j}$ $=\begin{cases}1&\text{if}\ j\leq i+\textit{W}\\ 0&\text{if}\ j>i+\textit{W}\end{cases}$ (4)
$M^{L}_{i,j}$ $=\prod_{n=j+1}^{i-1}(1-G_{n})$ (5)
where W is the sliding window size and $G_{n}$ a Global Token at Position $n$. We then construct the attention masks based on the type of each token. After that, we mask out the upper triangular part of the mask to ensure its causality.
In practice, we first collect the index of the next global token $I^{G_{\text{next}}}_{i}:=\min(\{k|k>=i\land G_{k}=1\})$ and rewrite Eq. 5 as:
$M^{L}_{i,j}=\begin{cases}1&\text{if}\ j\leq I^{G_{\text{next}}}_{i}\\ 0&\text{if}\ j>I^{G_{\text{next}}}_{i}\end{cases}$ (6)
During the backward process, we collect the column-wise gradient values from our attention masks and apply these gradients to update the parameters from our Attention Score Layer. Different from other approaches that approximates the attention maps or attention outputs from an existing model [8, 27, 85],  NAtS can be learned directly from the target labels, and thus, can be applied for both pre-training a new sparse transformer from scratch or fine tunning an existing transformer. To control the attention map’s sparsity, we introduce a small regularization value $\lambda$ that is directly applied to the gradient for Global Token and Local Tokens to encourage more Sliding Window Tokens. We show how the backward gradients is computed and how the $\lambda$ values control the attention map sparsity in the appendix  C.
These rules are then integrated in FlashAttention [17, 18] to avoid explicitly computing the attention masks during the forward pass. In addition to the transformer computation, we only need to collect the next Global Token indices $\mathbf{I^{G_{\text{next}}}}$ (with a complexity of $\mathcal{O}(N)$) and then mask out the attention map values with the masks defined above. At the same time, we also skip all the computation blocks that do not contain any valid values (i.e., all the items in the mask of that block are 0) during the online attention process. Further details can be found in the appendix D
3.3 Efficient Inference with Different Token Types
During inference time, we dynamically map the input feature maps to the corresponding token types and discard the tokens no longer required by the following tokens. The Sliding Window Tokens only survive for a fixed amount of time steps. We preserve a queue in the KV cache space that only stores the most recent W tokens and mask out the non-Sliding Window Tokens: when new tokens come, we place them after the tail of the queue to replace the tokens older than W.
Similar to the vanilla transformer forward process, when new tokens arrive, we concatenate them with the existing KV caches, generating new masks and computing the attention output. After that, we start our post-update process: we first check the state of each token to decide if we can remove them or keep them in the cached values. Since different heads might disagree on the types of the same token, we record both the sizes for Global Tokens ($Size_{G}$) and Local Tokens ($Size_{L}$) for all the heads. New Sliding Window Tokens do not change these sizes since they will always be placed separately. However, when a new Global Token for any head arrives, we remove all the Local Tokens from the corresponding heads and place the new Global Token right after the existing Global Tokens and then update our $Size_{G}$ and $Size_{L}$ accordingly. The same strategy is applied when new $LocalTokens$ arrive: we place them at the end of the Local Tokens and increase the number for Local Tokens.
Figure 5: An example of how caches are updated within NAtS when new tokens arrive with a model containing two heads. The two rows represent different heads.
Figure 5 illustrates an exemplary case to update the KV caches in NAtS with a sliding window size of 3. The first two places store the most recent tokens and use a mask to mask out the non-Sliding Window Tokens (yellow tokens). Since Token 8 for Head 1 is Sliding Window Token and Global Token for Head 2. We first move both tokens to the beginning of our cache to replace the old one. After that, we drop Token 8 in Head 1 since it has already moved to the sliding window caches. Then, since Token 8 in Head 2 is a Global Token, we drop all local tokens after the last Global Token (Token 1). Hence, we place Token 8 after Token 1 and remove Tokens 4 and 6. The $Size_{G}$ is thenupdated from $[5,3]$ to $[5,4]$ and the $Size_{L}$ is updated from $[0,2]$ to $[0,0]$: both new tokens are merged into the existing tokens, we do not need the extra space to store them.
4 Experiments
We implement NAtS based on the Flash Attention 2 [17] implementation on triton. In our experiments, we first train NAtS parameters jointly within a transformer model from scratch. We then apply NAtS to fine-tune a large language model to show that NAtS could efficiently reduce the KV cache size required while maintaining the model performance. The codes for  NAtS implementation and experimental designs can be found under: https://github.com/automl/NeuralAttentionSearch
4.1 Training a Transformer From Scratch
We first apply NAtS to train a GPT2 small style [67] transformer model with 128M parameters from scratch on the PG-19 Language Modeling Benchmark [68] with four Nvidia H100-PCIe GPUs and evaluate them on the test sets of PG19 with a context length of 1024. Further details can be found under the appendix  E.1.
As a baseline, we train another dense transformer model under the same hyperparameter setting. During inference time, we compare NAtS with the following baselines besides the full Transformer: (i) Streaming LLM [82] only preserves the first few starting and the most recent few tokens for future prediction. (ii) H2O [93] first computes the attention map and only preserves the tokens with the top-k attention scores. H2O and Streaming LLM are training-free approaches that control the sparsity with pre-defined hyperparameters during inference time. In contrast, NAtS controls this sparsity by the sparsity hyperparameter value $\lambda$. Hence, we train multiple models with different $\lambda$. However, in our experiments, we observe that the attention map sparsity values converge much faster than the model loss. We could quickly estimate the attention map sparsity within the first 10,000 iterations to check if this sparsity value satisfies the required sparsity and early-stop the runs that do not satisfy the requirements [38].
Figure 6: Perplexity vs KV Cache size under different sparsity settings $\lambda$ on the PG19 dataset.
We provide different hyperparameters for H2O and Streaming LLM for different sparsity (with the sliding window size of 32, 64, 128, 256 plus the same number of HH for H2O and 64, 128, 256, 512 plus 8 Sink tokens for Streaming LLM). This results in a corresponding KV budget fraction ranging from 6.25% to 50%. Meanwhile, we train NAtS with the following $\lambda$: 0,1e9,5e9,1e8,1e70,1e-9,5e-9,1e-8,1e-7.
Figure 6 shows the perplexity of the PG19 test sets from different model settings. The x-axis indicates the fraction of KV caches applied to generate the last token in the input sequence. As expected, larger $\lambda$ results in a smaller KV cache fraction. Interestingly, even if $\lambda$ is set as 0, the KV cache budget can still be reduced to around 35% with slightly higher perplexity compared to the full attention. This might indicate that the model also tends to sparsify the input information to focus more on the relevant parts from the input data [87].
H2O maintains nearly lossless perplexity with a budget of 50%. However, as the available budget decreases,  NAtS’s perplexity remains nearly the same until a KV cache size of around 10%. In contrast, H2O and Streaming LLM start to degenerate their performance starting from a $25\%$ budget allocation. The most sparse model in NAtS family only contains roughly $3\%$ of the KV cache size (meaning roughly 30 cached tokens per layer on average) and achieves a lower perplexity compared to H2O with $12.5\%$ of the KV caches.
4.2 Fine Tune a Large Language Model
We now apply NAtS to fine-tune an existing large language model [32, 40] in the long context scenario. We construct a new training dataset that follows the construction rules of LongBench as a training set to fine-tune LLM models with NAtS. Some of LongBench’s tasks are collected from the test sets of the previous benchmarks. Hence, we first collect the training sets from these benchmarks and construct these datasets following the data structure in LongBench. Additionally, we also add the passkey-retrieval dataset introduced in DuoAttention [81]. Overall, this dataset contains 7 0007\,000 instances with a maximal context length of 16 00016\,000. Further details can be found in Appendix E.2.
We only fine-tune the Attention Score Layer while keeping all the other parameters in the network fixed to approximate the original output from the corresponding base LLM. This approach is similar to DuoAttention [83]. However, since we want the model to capture the overall context information, we update Attention Score Layer with all the output from the full attention layer:
$\mathcal{L}_{distill}=\frac{1}{\mathit{B}}\sum_{i=1}^{\mathit{B}}\sum_{j=1}^{\mathit{L}}(\mathbf{H}_{full}^{(i)}[j]-\mathbf{H}_{NAtS}^{(i)}[j])^{2}$ (7)
where $\mathit{B}$ is the Batch Size. We only use the context information from the real-world dataset to ensure that the token information does not rely on specific prompts. However, the synthetic dataset might contain duplicated contexts, and hence, we only optimize the synthetic dataset with its corresponding labels [83]. Additionally, we control the sparsity with the regularization value $\lambda$ instead of the additional loss item defined in DuoAttention, and therefore, only optimize NAtS with $\mathcal{L}_{distill}$ as the loss function. We fine-tune two long-context models (Llama-3.1-8B-Instruct [32] and Mistral-7B-v0.3-Instruct [40]) on two Nvidia H100 PCIe GPUs for one epoch using AdamW [45, 57] with a learning rate of 2e-3 with a warm-up from 2e-4 in the first 20% steps and reducing back to 2e-4 in the last 20% steps. We apply different $\lambda$ to allow for different sparsity. For a fair comparison, we set the sliding window size W as 256, the same as DuoAttention.
In addition to the baselines in Section 4.1, we also evaluate other state-of-the-art KV eviction strategies, including SnapKV [52], AdaKV [25], ChunkKV [55], PyramidKV [14], CriticalKV [26]. These approaches rely on a sliding window from the query matrix during the pre-filling stage to identify the important tokens. Unlike NAtS and the other approaches listed above that update the KV caches on the fly, these approaches only evict the KV cache once after the prefilling stage. Additionally, we evaluate DuoAttention [83] and MoA [27]. Both approaches extend the existing streaming LLMs approach to provide different budgets for different attention heads. These two methods share the same idea as NAtS, where the importance of different KV caches should be learned instead of the pre-defined rules. However, instead of learning the importance of the head level, we aim at learning this information directly on the token level.
Since we only optimize the Attention Score Layer, the number of learnable parameters is $n_{layers}\cdot d_{model}\cdot n_{heads}\cdot{n_{options}}$, where $n_{options}$ is the number of options in our search space (in our case, it is $3$). Hence, the size of the parameters that need to be stored is negligible (i.e., it only takes roughly 1313MB on disk for each set of Attention Score Layer) compared to the weights of the LLM. Hence, users could pre-collect all these weights and apply the one that fits their compression requirement.
We evaluate all the baseline results on the RULER [36] and LongBench [11]. RULER is a synthetic benchmark that evaluates a model’s long-context capabilities across 4 task categories. LongBench evaluates the model’s capacity to understand the information in different long context scenarios. Following the setting from KV Press [63], we ask all the one-time KV eviction approaches to evict the KV cache once the model receives all the context information before the question-related information arrives. For the other approaches, we evict the KV caches dynamically when new tokens are generated.
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
4k 95.06 94.25 59.37 83.03 54.90 70.03 69.19 87.52 56.41 44.63 95.32 (51%)
8k 93.22 92.42 48.58 81.51 54.10 71.43 67.20 89.01 54.55 19.80 93.30 (45%)
16k 90.53 88.24 50.97 65.55 54.38 76.32 67.07 88.03 58.17 15.16 90.01 (42%)
32k 85.95 84.65 45.68 11.48 47.13 72.48 55.88 82.88 48.05 12.79 86.16 (41%)
64k 84.01 83.54 51.40 0.07 69.61 78.10 71.05 81.93 70.49 10.86 84.09 (41%)
128k 73.83 73.11 46.95 0.00 39.26 64.49 41.61 70.53 40.02 11.05 74.77 (44%)
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
4k 95.06 73.18 37.30 71.95 35.67 44.34 51.59 57.73 35.60 35.60 93.93 (24%)
8k 93.22 62.83 28.29 65.34 33.92 41.46 45.87 62.83 35.17 25.69 90.29 (19%)
16k 90.53 52.01 28.19 47.82 32.74 41.61 49.40 67.08 34.35 23.94 88.87 (16%)
32k 85.95 53.50 30.54 9.71 32.17 40.24 39.19 64.41 32.49 20.17 84.40 (15%)
64k 84.01 65.57 31.42 0.04 57.36 65.98 57.01 70.85 58.02 19.29 79.36 (15%)
128k 73.83 53.05 31.41 0.00 27.36 34.33 37.35 56.51 27.34 13.40 64.26 (16%)
Table 1: Ruler results with 50% KV budgets (top) and 25% KV budget (bottom) on LLama 3.1. All Full models use the 100% KV budgets. We mark the actual KV budgets used by NAtS in the bracket. Models with the best performance besides the base Full Attention Models are bold
The results on the Ruler Benchmark with LLama 3.1 are shown in Table  1. NAtS uses $\lambda=3e-7$ for the 25% KV size level and $\lambda=5e-8$ for the 50% KV size level. NAtS achieves nearly lossless scores across different input context lengths. When the KV budget drops to 25%, NAtS achieves 96% of the full attention scores with much smaller KV budget sizes, while the other baselines could no longer keep their performance with the decreased KV budgets. Results with Mistral-7B-Instruct-v0.3 can be found under Appendix F.
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
NarrativeQA 31.35 22.62 27.83 21.07 27.08 29.02 26.27 30.48 28.02 19.19 30.53(13%)
Qasper 24.73 18.99 14.53 15.00 14.24 14.84 12.71 15.79 13.43 18.57 23.76(27%)
MultiFieldQA-en 29.46 25.36 16.47 18.24 17.72 19.31 17.44 21.74 17.72 14.75 28.94(24%)
MultiFieldQA-zh 60.01 54.95 33.94 29.14 33.77 35.34 33.49 36.30 32.76 21.23 61.18(22%)
HotpotQA 17.06 16.06 13.95 16.47 15.58 16.12 16.50 16.73 14.23 10.18 16.06(20%)
2WikiQA 16.64 16.18 13.53 11.28 13.45 13.17 13.41 13.89 12.36 11.42 17.05(23%)
Musique 11.59 8.41 8.96 11.13 10.81 10.68 11.05 11.25 9.40 5.93 11.42(18%)
DuReader (zh) 35.56 32.18 18.64 24.77 23.57 24.57 24.24 24.68 23.56 22.33 34.28(17%)
GovReport 34.30 27.63 28.20 26.90 28.51 28.28 28.86 29.58 27.53 25.24 33.82(19%)
QMSum 23.30 22.28 20.67 20.39 21.41 22.05 21.27 22.27 21.98 19.59 23.11(17%)
MultiNews 27.13 24.58 21.76 22.80 23.14 23.50 22.72 24.03 23.00 23.30 26.72(33%)
VCSUM (zh) 16.36 14.97 14.26 14.85 14.80 15.18 14.92 15.46 15.21 14.58 15.61(14%)
TREC 72.50 58.50 66.00 55.00 53.50 59.00 54.50 62.00 48.50 60.50 72.00(26%)
TriviaQA 91.15 82.81 90.69 89.98 91.00 91.55 90.72 90.97 90.92 71.51 91.61(22%)
SAMSum 43.72 40.14 41.97 41.98 43.59 43.11 42.32 43.79 43.15 42.30 44.21(16%)
LSHT 46.50 35.00 35.50 27.50 45.00 44.00 43.50 46.00 45.00 22.00 47.50(16%)
Passage Count 6.63 4.50 7.22 3.40 5.52 5.14 4.11 5.90 4.70 3.79 7.12(20%)
PassageRetrieval-en 97.98 91.65 94.81 72.25 90.36 90.70 90.38 94.70 88.32 28.21 95.81(19%)
PassageRetrieval-zh 77.99 55.22 28.97 52.13 73.06 78.74 55.95 80.15 74.66 24.78 78.5(20%)
LCC 54.10 54.27 52.38 54.52 55.40 54.93 54.12 54.26 54.90 54.14 53.54(38%)
RepoBench-P 51.39 57.26 49.65 55.54 51.80 52.44 52.87 52.57 52.41 53.35 53.48(28%)
Table 2: LongBench Results with 25% Budget Allocation. The Full model uses the 100% KV budgets. The numbers in the brackets for NAtS is the used value KV cache sizes. We bold the models with the best performance besides the base Full Attention Models .
For the LongBench tasks, we evaluate all the baselines with $50\%$ and $25\%$ of the KV cache sizes. We show the result with KV cache size of $25\%$ for LLama3.1-8B in Table 2. The model that is used in this task is a model with $\lambda=3e-7$. NAtS achieves the best performance on most of the datasets with (in many cases) smaller KV cache values. We provide further results, including results with Mistral-7B and  NAtS trained with other parameters, in Appendix  F.
4.3 Latency Evaluation
[…]
Figure 7: Memory and latency usage during pre-filling (left) and decoding (right)
5 Conclusion and Future Work
Efficiently managing the KV cache is crucial for deploying large language models in long-context applications. In this work, we propose NAtS, an approach that automatically optimizes the optimal roles of each token to determine how long they should survive. By constructing a learnable mask, NAtS learns to sparsify the attention map end-to-end. Our experiments show that NAtS uses much less KV caches compared to the State-of-the-art KV caching reduction approach. While NAtS shows promising results, future work could include exploration of further token roles and structured search spaces with hierarchies. Overall, we believe that NAtS paves the way towards more efficient and scalable inference with LLMs.
6 Acknowledgement
The authors gratefully acknowledge the computing time provided to them on the high-performance computers Noctua2 at the NHR Center PC2 under the project hpc-prf-intexml. These are funded by the Federal Ministry of Education and Research and the state governments participating on the basis of the resolutions of the GWK for the national high performance computing at universities (www.nhr-verein.de/unsere-partner).
Difan Deng was supported by the Federal Ministry of Education and Research (BMBF) under the project AI service center KISSKI (grantno.01IS22093C).
Appendix A Limitations
While NAtS shows promising results that efficiently reduce the KV cache sizes and our search space already contains most of the KV cache eviction strategies. However, the construction of the attention mask is still column-wise oriented and mostly focuses on the vertical style sparse attention [41]. Hence, this work does not involve the block-sparse attention and slash attention that are widely observed in other frameworks [41, 48]. One potential future direction would be to introduce a more structured search space that allows for the construction of more sparse attention types.
Appendix B Broader Impact
LLMs are widely applied to different fields nowadays. However, the cost for LLM to store the KV cache and predict the next token is still huge, given the $\mathcal{O}(\mathit{L})$ computation and memory costs of full Attention Models. This prevents further adaptation of LLMs (and other transformer-based foundation models with properties similar to those described before) because of high energy consumption and limited context windows. NAtS achieves a substantial reduction in KV cache size with minimal impact on model performance, outperforming existing state-of-the-art approaches. This increased efficiency can enable the deployment of larger, more powerful language models on resource-constrained devices and facilitate the development of new applications that rely on long-context understanding, such as advanced conversational AI, comprehensive document summarization, and complex code generation. By making long-context processing more accessible, NAtS has the potential to accelerate progress in natural language processing and related fields. Nevertheless, it does not solve other inherent problems of LLMs such as hallucinations.
Appendix C Details on Backward Propagation
C.1 Gradients for Attention Masks
To compute the gradients for M, we set $g(A,M)=e^{A}\odot M$; then the gradient for M is:
$\frac{\partial O}{\partial M}$ $=\frac{\partial O}{\partial g}\frac{\partial g}{\partial M}$ (8)
$\frac{\partial g}{\partial M}$ $=e^{A}$ (9)
$\frac{\partial g}{\partial A}$ $=e^{A}\odot M$ (10)
In Eq. 9 and  10, we show the gradient for $M$ is the same as the value that $\partial O/\partial A$ is supposed to be if no mask is applied. Since we have $g_{i,j}=e^{A_{i,j}}\odot M_{i,j}$. Let’s set $S_{i}:=\sum_{j}g_{i,j}$ and $P_{i,j}:=\frac{g_{i,j}}{S_{i}}$. Then we have $\mathbf{d}\mathbf{P}=\mathbf{d}\mathbf{O}\mathbf{V}^{T}$. Therefore,
$dg_{i:}=(diag(\frac{1}{S_{i:}})-\frac{1}{S_{i:}}P_{i:}^{T})dP_{i:}$ (11)
Combining Eq. 11 with Eq. 10 and Eq. 9 provides the same gradients as the vanilla softmax function with additive masks:
$dM_{i:}$ $=(diag(\frac{e^{A_{i:}}}{S_{i:}})-\frac{e^{A_{i:}}}{S_{i:}}P_{i:}^{T})dP_{i:}$ (12)
$dA$ $=dM\odot M$
$=(diag(P_{i:})-P_{i:}P^{T}_{i:})dP_{i:}$ (13)
Since $dg_{i:}$ is required to compute the gradient for $\mathbf{A}_{i:}$ and always needs to be computed. We can directly use this information to compute the gradients for the attention Mask $M$.
Following FlashAttention [18], we define $D_{i}=do_{i}^{T}o_{i}$, then
$dM_{i,j}=\frac{e^{A_{i,j}}}{S_{i:}}(dP_{ij}-D_{i})$~ (14)
$dA_{i,j}=P_{i,j}(dP_{ij}-D_{i})$~ (15)
and $dA$ is computed by Eq. 13. After that, we can backpropagate $dA$ to $dq$ and $dv$. Since $M$ needs to be recomputed anyway in the flash attention’s backward process, this only results in little computational overhead.
However, in practice, we cannot guarantee an upper bound for $\frac{e^{A_{i,j}}}{S_{i}}$ with $M_{i,j}=0$ since $S_{i}:=\sum_{j}e^{A_{i,j}}\odot M_{i,j}$. As a result, $\lim_{e^{(A_{i,j})}\to\infty}(\frac{e^{(A_{i,j})}}{S_{i}})=\infty$ when $M_{i,j}=0$. Hence, we first clip $\frac{e^{A_{i,j}}}{S_{i}}$ within $(0,1)$ and then compute $dM$ with the clipped value:
$P^{\prime}_{i,j}$ $=\min(\frac{e^{A_{i,j}}}{S_{i}},1)$ (16)
$dM_{i,j}$ $=P^{\prime}_{i,j}(dP_{ij}-D_{i})$~ (17)
C.2 Details on computing the gradients for computing the token gradients info $d\alpha$
The gradients towards each token are collected through the column sum of each value weighted by the corresponding attention masks:
$d\alpha_{i}=\sum_{j}dM_{i,j}\times M_{i,j}^{\alpha}$~ (18)
Where $\alpha\in\{G,L,SW\}$ is the discretized token type. Intuitively, this shows the model’s preference over short-range or long-range correlations: If $i$ is quite close to $j$, then all the $\alpha$ will receive the same gradient information. However, if the model wants to create a long-range correlation with $i\gg j$, only Global Tokens will receive the gradient information. The network will therefore prefer to classify the corresponding tokens as global tokens.
Eq. 5 shows that the Global Token controls the local mask size. Therefore, the gradients for Global Token $i$ should also be regularized by the gradient information from the previous tokens:
$\frac{\partial M_{i,j}^{L}}{\partial G_{k}}=-\prod_{\begin{subarray}{c}n=j+1\\ n\neq k\end{subarray}}^{i-1}(1-G_{n})$~ (19)
In cases where $G_{k}$ is 0, Equation 19 is the negative value of Equation 5. However, for the case where $G_{k}=1$, this is equivalent to the local mask values where $k$ is no longer set as a Global Token. This requires us to find the index of the next global token $I^{G_{\text{next}}}_{k+1}$ and the last global token $I^{G_{\text{next}}}_{k-l}$ where $l$ is the length of the local sequence that ends at $k$.
This gradient information will then be collected and subtracted from the computed Global Tokens gradient values:
$d\alpha^{G-}_{i}$ $=\begin{cases}\sum\limits_{\begin{subarray}{c}m\geq i\\ i>n\geq I^{G_{\text{next}}}_{k-l}\end{subarray}}M^{L}_{m,n}\times dM_{m,n}\ \ &\text{if}\ G_{i}=0,\\ \sum\limits_{\begin{subarray}{c}I^{G_{\text{next}}}_{i+1}\geq m>=i\\ i>n\geq I^{G_{\text{next}}}_{k-l}\end{subarray}}dM_{m,n}\ \ &\text{if}\ G_{i}=1,\end{cases}$ (20)
$d\alpha^{G}_{i}$ $=d\alpha^{G}_{i}-d\alpha^{G-}_{i}$ (21)
Intuitively, this gradient term $d\alpha^{G-}_{i}$ checks if the new Global Token needs to be inserted into the current sub-sequence (when $G$ is 0) or we should remove the current Global Token to enlarge the current sub-sequence (when $G$ is 1). We further illustrate this process in the appendix.
Figure  8 illustrates an example of this. Token 4 is a global token, and we search for its next Global Token (which is Token 10 in this example). Assuming that we want to change Token 4 to another role, the regions within the boundary (orange ones) are those tokens that are influenced by this swtching: given that Token 4 no longer becomes Global Token, Token 2 and 3 will be part of a larger subsequence that ranges from 1 to 10, and the attentino maps within the orange region should not be masked out. Intuitively, this gradient measures the regret that we made in order to switch one Global Token into a Local Token. A similar idea can be found in the red region. Assuming that we want to switch Token 7 to a Global Token, then Tokens 5,6 will be split into a new subsequence and we will no longer connect them with Tokens 8, 9, 10 since they belong to two different sub-sequences after the switch. Hence, this value in the red regions measures the regret if we mistakenly classify a Global Token as non-Global Tokens.
Figure 8: The gradient term $d\alpha^{G-}_{i}$ for Token 4 and 7
In practice, the $d\alpha^{G-}_{i}$ values for Global Token can be easily collected by looking at the last and next $G$ index with a scan function to collect the gradients to the corresponding positions. However, computing $d\alpha^{G-}_{i}$ with $G_{i}=0$ requires the gradient information from all the past token mask information. This might be computationally prohibitive in practice, since the valid values contained in Equation 20 are sparse, we ignore the cases for $G_{i}=0$ and only compete the dense gradeints for the case where $G_{i}=1$.
C.3 Optimizing for sparser Attention Maps
The sparse regularization term $\lambda$ is directly applied to the corresponding gradients for the Global Tokens and Local Tokens. This value should penalize the number of unmasked tokens for each row. While the number of unmasked tokens for Global Token and Local Tokens in column $i$ are $\mathit{L}-i$ and $I^{G_{\text{next}}}_{i}-i$ with $I^{G_{\text{next}}}_{i}$ defined in Section 3.2. Hence, we have:
$dG^{G_{sparse}}_{i}$ $=\lambda\times\frac{\mathit{L}-i}{\mathit{L}}$ (22)
$dG^{L_{sparse}}_{i}$ $=\lambda\times\frac{I^{G_{\text{next}}}_{i}-i}{\mathit{L}}$ (23)
Combining Eq.  18,  20,  22 and Eq.  23, we have:
$d\alpha^{G}_{i}$ $=\sum_{j}dM_{i,j}\times M_{i,j}^{G}\times M_{i,j}^{casual}+dG^{G_{sparse}}_{i}-d\alpha^{G-}_{i}$ (24)
$d\alpha^{L}_{i}$ $=\sum_{j}dM_{i,j}\times M_{i,j}^{L}\times M_{i,j}^{casual}+dG^{L_{sparse}}_{i}$ (25)
$d\alpha^{SW}_{i}$ $=\sum_{j}dM_{i,j}\times M_{i,j}^{SW}\times M_{i,j}^{casual}$ (26)
with $M_{i,j}^{G},M_{i,j}^{SW},M_{i,j}^{L}$ defined in Eq.  3,  6 and  4 and $M_{i,j}^{casual}$ is a casual attention mask.
Appendix D Integrating NAtS into FlashAttention
Algorithm 1 NAtS forward pass on FlashAttention2, we mark the  NAtS related operations with blue
0: Matrices $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d}$, token states $\mathbf{\alpha}\in\mathbb{R}^{N\times\mathit{N_{\text{opts}}}}$ , indices of the next token states $\mathbf{I^{G_{\text{next}}}}\in\mathbb{R}^{N}$ in HBM, block sizes $B_{c}$, $B_{r}$, sliding window size W
1: Divide $\mathbf{Q}$ into $T_{r}=\left\lceil\frac{N}{B_{r}}\right\rceil$ blocks $\mathbf{Q}_{1},\dots,\mathbf{Q}_{T_{r}}$ of size $B_{r}\times d$ each, and divide $\mathbf{K},\mathbf{V}$, $\mathbf{\alpha},\mathbf{I^{G_{\text{next}}}}$ into $T_{c}=\left\lceil\frac{N}{B_{c}}\right\rceil$ blocks $\mathbf{K}_{1},\dots,\mathbf{K}_{T_{c}}$, $\mathbf{V}_{1},\dots,\mathbf{V}_{T_{c}}$, of size $B_{c}\times d$, $\mathbf{\alpha}_{1},\dots,\mathbf{\alpha}_{T_{c}}$ of size $B_{r}\times\mathit{N_{\text{opts}}}$ , and $\mathbf{I^{G_{\text{next}}}}_{1},\dots,\mathbf{I^{G_{\text{next}}}_{T_{c}}}$ of size $B_{r}$ each.
2: Divide the output $\mathbf{O}\in\mathbb{R}^{N\times d}$ into $T_{r}$ blocks $\mathbf{O}_{i},\dots,\mathbf{O}_{T_{r}}$ of size $B_{r}\times d$ each, and divide the logsumexp $L$ into $T_{r}$ blocks $L_{i},\dots,L_{T_{r}}$ of size $B_{r}$ each.
3:for $1\leq i\leq T_{r}$ do
4:   Load $\mathbf{Q}_{i}$ from HBM to on-chip SRAM.
5:   On chip, initialize $\mathbf{O}_{i}^{(0)}=(0)_{B_{r}\times d}\in\mathbb{R}^{B_{r}\times d},\ell_{i}^{(0)}=(0)_{B_{r}}\in\mathbb{R}^{B_{r}},m_{i}^{(0)}=(-\infty)_{B_{r}}\in\mathbb{R}^{B_{r}}$.
6:  for $1\leq j\leq T_{c}$ do
7:   Load $\mathbf{\alpha}_{j}$ and $\mathbf{I^{G_{\text{next}}}}_{j}$ to SRAM and check if the current block contain any valid values according to Equations 3,4,6,
8:   if Do Compute then
9:    Construct attention mask $\mathbf{M}_{i,j}$ with $\mathbf{\alpha}_{j}$, $\mathbf{I^{G_{\text{next}}}}_{j}$, and W with Equations 3,4,6
10:     Load $\mathbf{K}_{j},\mathbf{V}_{j}$ from HBM to on-chip SRAM.
11:     On chip, compute $\mathbf{S}_{i}^{(j)}=\mathbf{Q}_{i}\mathbf{K}_{j}^{T}\in\mathbb{R}^{B_{r}\times B_{c}}$.
12:     On chip, compute $m_{i}^{(j)}=\mathrm{max}(m_{i}^{(j-1)},\mathrm{rowmax}(\mathbf{S}_{i}^{(j)}))\in\mathbb{R}^{B_{r}}$, $\tilde{\mathbf{P}}_{i}^{(j)}=\exp(\mathbf{S}_{i}^{(j)}-m_{i}^{(j)}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\odot\mathbf{M}_{i,j}}\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\ell_{i}^{(j)}=e^{m_{i}^{j-1}-m_{i}^{(j)}}\ell_{i}^{(j-1)}+\mathrm{rowsum}(\tilde{\mathbf{P}}_{i}^{(j)})\in\mathbb{R}^{B_{r}}$.
13:     On chip, compute $\mathbf{O}_{i}^{(j)}=\mathrm{diag}(e^{m_{i}^{(j-1)}-m_{i}^{(j)}})^{-1}\mathbf{O}_{i}^{(j-1)}+\tilde{\mathbf{P}}_{i}^{(j)}\mathbf{V}_{j}$.
14:   end if
15:  end for
16:  On chip, compute $\mathbf{O}_{i}=\mathrm{diag}(\ell_{i}^{(T_{c})})^{-1}\mathbf{O}_{i}^{(T_{c})}$.
17:  On chip, compute $L_{i}=m_{i}^{(T_{c})}+\log(\ell_{i}^{(T_{c})})$.
18:  Write $\mathbf{O}_{i}$ to HBM as the $i$-th block of $\mathbf{O}$.
19:  Write $L_{i}$ to HBM as the $i$-th block of $L$.
20:end for
21: Return the output $\mathbf{O}$ and the logsumexp $L$.
Flashattention [17] continuously loads Q, K, V values to the SRAM and only computes the attention maps within the loaded blocks.  NAtS loads the corresponding token states information $\alpha$ and construct the masks $M$ dynamically with Equation 3,  5,  4. During the backward process, Flashattention collects the row-wise gradients for computing $d\mathbf{Q}$ and column-wise gradients to compute $d\mathbf{K}$ and $d\mathbf{V}$. Hence, the gradients $d\mathbf{\alpha}$ can be computed together with $d\mathbf{K}$ and $d\mathbf{V}$. We show how  NAtS can be integrated into Flashattion2  [17] in Algorithms 1 and 2. We highlight the difference between the vanilla Flashattention2 and Flashattention2 with  NAtS in blue. Compared with vanilla Flashattention models,  NAtS adaptively skips the blocks that do not have valid mask values to avoid unnecessary computation. These invalid blocks will only be applied to update the gradients for $d\mathbf{\alpha}$.
Algorithm 2 NAtS backward pass on FlashAttention2, we mark the  NAtS related operations with blue
0: Matrices $\mathbf{Q},\mathbf{K},\mathbf{V},\mathbf{O},\mathbf{dO}\in\mathbb{R}^{N\times d}$, token states $\mathbf{\alpha}\in\mathbb{R}^{\mathit{L}\times\mathit{N_{\text{opts}}}}$ , indices of the next token states $\mathbf{I^{G_{\text{next}}}}\in\mathbb{R}^{\mathit{L}}$ in HBM, vector $L\in\mathbb{R}^{N}$ in HBM, block sizes $B_{c}$, $B_{r}$, sliding window size W.
1: Divide $\mathbf{Q}$ into $T_{r}=\left\lceil\frac{N}{B_{r}}\right\rceil$ blocks $\mathbf{Q}_{1},\dots,\mathbf{Q}_{T_{r}}$ of size $B_{r}\times d$ each, and divide $\mathbf{K},\mathbf{V}$, $\mathbf{\alpha},\mathbf{I^{G_{\text{next}}}}$ in to $T_{c}=\left\lceil\frac{N}{B_{c}}\right\rceil$ blocks $\mathbf{K}_{1},\dots,\mathbf{K}_{T_{c}}$ and $\mathbf{V}_{1},\dots,\mathbf{V}_{T_{c}}$, of size $B_{c}\times d$, $\mathbf{\alpha}_{1},\dots,\mathbf{\alpha}_{T_{c}}$ of size $B_{r}\times\mathit{N_{\text{opts}}}$ , and $\mathbf{I^{G_{\text{next}}}}_{1},\dots,\mathbf{I^{G_{\text{next}}}_{T_{c}}}$ of size $B_{r}$ each.
2: Divide $\mathbf{O}$ into $T_{r}$ blocks $\mathbf{O}_{i},\dots,\mathbf{O}_{T_{r}}$ of size $B_{r}\times d$ each, divide $\mathbf{dO}$ into $T_{r}$ blocks $\mathbf{dO}_{i},\dots,\mathbf{dO}_{T_{r}}$ of size $B_{r}\times d$ each, and divide $L$ into $T_{r}$ blocks $L_{i},\dots,L_{T_{r}}$ of size $B_{r}$ each.
3: Initialize $\mathbf{dQ}=(0)_{N\times d}$ in HBM and divide it into $T_{r}$ blocks $\mathbf{dQ}_{1},\dots,\mathbf{dQ}_{T_{r}}$ of size $B_{r}\times d$ each. Divide $\mathbf{dK},\mathbf{dV}\in\mathbb{R}^{N\times d}$ in to $T_{c}$ blocks $\mathbf{dK}_{1},\dots,\mathbf{dK}_{T_{c}}$ and $\mathbf{dV}_{1},\dots,\mathbf{dV}_{T_{c}}$, of size $B_{c}\times d$ each. Divide $d\mathbf{\alpha}\in\mathbb{R}^{\mathit{N_{\text{opts}}}}$ in to $d\mathbf{\alpha}_{1},\dots,d\mathbf{\alpha}_{T_{c}}$ of size $B_{r}\times\mathit{N_{\text{opts}}}$ each.
4: Compute $D=\mathrm{rowsum}(\mathbf{dO}\odot\mathbf{O})\in\mathbb{R}^{d}$ (pointwise multiply), write $D$ to HBM and divide it into $T_{r}$ blocks $D_{1},\dots,D_{T_{r}}$ of size $B_{r}$ each.
5:for $1\leq j\leq T_{c}$ do
6:  Load $\mathbf{K}_{j},\mathbf{V}_{j}$, $\mathbf{\alpha}_{j}$ and $\mathbf{I^{G_{\text{next}}}}_{j}$ from HBM to on-chip SRAM.
7:  Initialize $\mathbf{dK}_{j}=(0)_{B_{c}\times d},\mathbf{dV}_{j}=(0)_{B_{c}\times d}$, $d\mathbf{\alpha}_{j}=(0)_{B_{c}\times\mathit{N_{\text{opts}}}}$ on SRAM.
8:  for $1\leq i\leq T_{r}$ do
9:   Load $\mathbf{Q}_{i},\mathbf{O}_{i},\mathbf{dO}_{i},\mathbf{dQ}_{i},L_{i},D_{i}$ from HBM to on-chip SRAM.
10:   On chip, compute $\mathbf{S}_{i}^{(j)}=\mathbf{Q}_{i}\mathbf{K}_{j}^{T}\in\mathbb{R}^{B_{r}\times B_{c}}$.
11:   On chip, Construct attention mask $\mathbf{M}_{i,j}$ with $\mathbf{\alpha}_{j}$, $\mathbf{I^{G_{\text{next}}}}_{j}$, and W with Equations 3,4,6
12:   On chip, compute $\mathbf{P^{\prime}}_{i}^{(j)}=\exp(\mathbf{S}_{ij}-L_{i})\in\mathbb{R}^{B_{r}\times B_{c}}$ and $\mathbf{P}_{i}^{(j)}=\mathbf{P^{\prime}}_{i}^{(j)}\odot\mathbf{M}_{i,j}$.
13:   Check if the current block contain any valid value
14:   if Do compute then
15:    On chip, compute $\mathbf{dV}_{j}\leftarrow\mathbf{dV}_{j}+(\mathbf{P}_{i}^{(j)})^{\top}\mathbf{dO}_{i}\in\mathbb{R}^{B_{c}\times d}$.
16:    On chip, compute $\mathbf{dP}_{i}^{(j)}=\mathbf{dO}_{i}\mathbf{V}_{j}^{\top}\in\mathbb{R}^{B_{r}\times B_{c}}$.
17:    On chip, compute $\mathbf{dS}_{i}^{(j)}=\mathbf{P}_{i}^{(j)}\odot(\mathbf{dP}_{i}^{(j)}-D_{i})\in\mathbb{R}^{B_{r}\times B_{c}}$.
18:    Load $\mathbf{dQ}_{i}$ from HBM to SRAM, then on chip, update $\mathbf{dQ}_{i}\leftarrow\mathbf{dQ}_{i}+\mathbf{dS}_{i}^{(j)}\mathbf{K}_{j}\in\mathbb{R}^{B_{r}\times d}$, and write back to HBM.
19:    On chip, compute $\mathbf{dK}_{j}\leftarrow\mathbf{dK}_{j}+{\mathbf{dS}_{i}^{(j)}}^{\top}\mathbf{Q}_{i}\in\mathbb{R}^{B_{c}\times d}$.
20:   end if
21:   On chip, clip $\mathbf{P^{\prime}}_{i}^{(j)}$ to (0,1) and compute $d\mathbf{S^{\prime}}_{i}^{(j)}=\mathbf{P^{\prime}}_{i}^{(j)}\odot(\mathbf{dP}_{i}^{(j)}-D_{i})\in\mathbb{R}^{B_{r}\times B_{c}}$
22:   On chip, update $d\mathbf{\alpha}_{j}$ with Equations 24, 25, 26
23:  end for
24:  Write $\mathbf{dK}_{j},\mathbf{dV}_{j}$, and $d\mathbf{\alpha}_{j}$ to HBM.
25:end for
26: Return $\mathbf{dQ},\mathbf{dK},\mathbf{dV}$.
Appendix E Experiments Details
Here we discuss further details in our experiments. The codes for  NAtS can be found under  https://github.com/automl/NeuralAttentionSearch
E.1 Detailed Hyperparameter Setting for GPT2-small Training
Following the setting from NanoGPT [43], the GPT-2 style small has 12 layers and 12 heads with a hidden dimension of 768. Instead of the learnable position encoding, we apply rotary embeddings [72] to each transformer layer. The PG-19 dataset contains books extracted from Project Gutenberg [65] with about 2B tokens in the training sets. We train all models with a context length of 1024 and a batch size of 480 (using gradient accumulation). We train them for 600 000600\,000 iterations evaluate them on the test sets of PG19 with a context length of 1024. Training one model can be finished in 16 hours with 4 Nvidia H100 PCEi GPUs.
E.2 Collecting the Fine-tune Training Set
To fine-tune  NAtS on LLMs, we collect the training datasets from different tasks for real-world tasks:
  • Multi-Document QA: HotPotQA [86], 2WikiMultihopQA [35], MuSiQue [76], and DuReader (zh) [34]
  • Single-Document QA: NarrativeQA [46] and Qasper [19]
  • Summarization: GovReport [37], QMSum [94], MultiNews [23], and VCSUM (zh) [80]
  • Few-shot Leraning: TREC [51], TriviaQA [42], and SAMSum [31]
  • Code Completion: LCC [33] and RepoBench-P [54]
We also construct the synthetic passkey-retrieval dataset introduced in DuoAttention [83]. This dataset is generated by embedding multiple random passkeys in different locations with a long context. The model will then be asked to recall this passkey information.
For all the few-shot learning datasets, following  Bai et al. [11], we randomly concatenate multiple question-answer pairs into one single extended context as one piece of data. The number of concatenated samples for the TREC dataset ranges from $[10,100]$. This value is $[2,6]$ for TriviaQA and $[10,50]$ for SAMSum. Additionally, for the datasets that do not have enough context length (e.g., the DuReader dataset), we also merge multiple documents into one piece of data (in our case, this value is 4).
We do not collect all the data whose length goes beyond a threshold to ensure that the context can be fitted into our model. Additionally, we collect at most 500 instances in each dataset since some datasets might not contain data. In the end, the real-world dataset contains 6436 data instances. We add another 564 instances from the synthetic dataset. In the end, our dataset contains 7000 instances. Finetuning a model on this dataset for one epoch takes roughly 8 hours with 2 Nvidia H100 PCIe GPUs.
Appendix F Further Experimental Results
F.1 Results on Ruler Benchmark with Mistral
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
4k 93.68 92.63 58.89 5.60 46.50 57.92 59.33 74.06 44.00 55.73 93.68 (49%)
8k 91.29 89.94 59.21 2.82 38.12 49.30 53.53 83.12 37.66 41.98 90.76 (43%)
16k 89.85 87.51 55.18 3.00 38.59 52.80 63.46 82.83 38.05 34.31 89.09 (41%)
32k 81.24 78.81 47.63 2.70 49.32 66.20 71.91 76.23 46.16 27.92 80.39 (42%)
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
4k 93.68 52.44 42.96 3.26 33.72 38.79 42.74 46.11 33.66 33.26 93.18 (29%)
8k 91.29 40.96 36.54 2.37 31.36 36.67 42.01 46.73 28.64 25.47 89.97 (24%)
16k 89.85 35.42 33.38 1.92 30.30 37.48 42.84 57.17 28.17 25.55 87.69 (21%)
32k 81.24 40.00 32.76 1.70 35.17 44.46 60.56 61.55 31.84 24.88 78.60 (21%)
Table 3: Ruler results with 50% KV budgets (top) and 25% KV budget (bottom) on Mistral-7B-Instruct-v0.3. All Full models use the 100% KV budgets. We mark the actual KV budgets used by  NAtS in the bracket.
Table  3 shows the results of different optimizers on the Mistral-7B-Instruct-v0.3 model. In this case, the $\lambda$ for 25% and 50% budgets are 1e61e-6 and 3e73e-7 respectively. We found that to achieve the same sparsity level, Mistral-7B requires a larger sparse regularization value $\lambda$ compared to the Llama 3.1 model. This might indicate that the attention map values in Mistral are more evenly distributed, and thus we need a larger $\lambda$ to force more tokens to only focus on the local information. The maximal context length of Mistral 7B is $32k$. We only evaluate the models with a maximal length of $32k$. The result is consistent with the results shown in Table 1.  NAtS consistently outperforms the other baselines with smaller budgets.
F.2 Further results on LongBench
We show additional results on fine-tuning LLM on the LongBench dataset here. Table 4 shows the results with 25% KV budgets on Mistral-7B-Instruct-v0.3 with $\lambda=3e-6$. The result confirms our conclusion that  NAtS outperforms the other baselines in most datasets under a similar budget level (and many times with even smaller budgets).
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
NarrativeQA 29.21 16.49 27.56 23.65 22.70 24.34 21.22 27.72 21.97 21.67 26.7(9%)
Qasper 41.34 14.58 29.16 18.95 23.08 22.99 21.21 28.85 23.54 19.72 41.46(21%)
MultiFieldQA-en 52.51 29.83 31.88 31.52 38.56 38.11 34.11 45.18 35.34 28.80 52.36(18%)
MultiFieldQA-zh 58.03 29.15 30.47 29.21 31.42 32.30 32.40 38.12 29.67 23.20 54.13(16%)
HotpotQA 49.62 34.85 46.07 42.10 46.37 48.83 46.93 47.48 45.94 39.64 51.45(15%)
2WikiQA 40.01 27.83 35.11 30.43 33.22 33.14 32.03 38.24 33.96 27.02 37.05(18%)
Musique 28.44 13.03 22.65 16.99 22.45 21.28 22.62 25.01 23.67 16.22 27.88(14%)
DuReader (zh) 34.87 26.51 18.69 20.66 25.34 27.55 28.45 27.04 24.82 19.11 35.25(12%)
GovReport 34.94 22.09 27.44 27.11 28.31 27.94 28.94 29.33 26.94 25.58 33.39(14%)
QMSum 25.67 17.25 22.11 20.84 21.43 21.95 21.95 22.77 22.39 19.53 24.70(13%)
MultiNews 27.85 23.41 22.01 23.03 23.77 23.62 23.34 24.37 23.89 22.55 27.35(28%)
VCSUM (zh) 16.34 13.58 15.41 15.13 15.28 15.38 14.69 16.10 14.96 14.87 16.06(11%)
TREC 75.50 52.00 70.50 61.50 58.50 59.00 57.00 67.50 58.00 63.00 73.50(20%)
TriviaQA 88.89 84.85 88.92 86.60 88.89 89.39 89.19 88.98 89.31 86.94 89.80(17%)
SAMSum 47.29 42.15 45.44 45.16 46.40 46.72 46.51 46.67 46.58 43.90 46.91(13%)
LSHT 39.75 17.50 29.00 22.00 38.50 39.00 39.50 39.25 37.50 17.25 38.5(11%)
Passage Count 5.50 3.50 6.00 6.00 6.00 8.00 5.50 6.00 6.50 5.00 4.5(14%)
PassageRetrieval-en 98.00 62.50 86.50 63.00 91.50 94.00 87.50 89.50 88.00 17.25 96.00(13%)
PassageRetrieval-zh 96.50 10.50 26.50 34.50 69.25 83.25 71.00 90.00 75.50 16.00 92.00(14%)
LCC 53.02 50.87 51.22 52.77 55.67 55.67 53.76 56.17 55.46 51.08 53.81(29%)
RepoBench-P 56.83 48.93 54.80 55.74 56.82 57.06 56.75 56.75 56.40 53.59 56.7(23%)
Table 4: LongBench Results with 25% Budget Size on Mistral-7B. .
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
NarrativeQA 31.35 29.40 28.18 20.15 30.37 30.04 29.49 32.03 29.18 13.92 31.39(29%)
Qasper 24.73 20.37 18.15 18.09 18.44 19.42 17.24 22.13 18.54 22.25 25.17(45%)
MultiFieldQA-en 29.46 26.70 17.89 22.50 23.21 24.62 22.55 26.30 22.95 14.24 28.94(41%)
MultiFieldQA-zh 60.01 60.31 42.19 38.88 47.66 50.23 44.54 56.22 46.50 23.91 61.27(37%)
HotpotQA 17.06 18.84 16.73 15.61 16.55 16.29 17.22 15.83 15.86 8.85 17.28(37%)
2WikiQA 16.64 16.30 14.28 11.71 15.82 15.38 14.67 17.01 14.81 10.58 16.84(41%)
Musique 11.59 13.77 10.75 10.12 10.78 11.19 10.63 12.85 10.64 5.03 11.89(36%)
DuReader (zh) 35.56 33.29 18.16 27.32 29.22 30.82 28.49 32.81 27.93 22.80 33.36(32%)
GovReport 34.30 32.99 30.84 29.46 31.34 31.51 31.92 33.09 30.54 25.45 34.23(36%)
QMSum 23.30 23.89 21.43 20.72 22.08 22.74 22.22 22.89 22.92 19.76 22.6(33%)
MultiNews 27.13 26.29 24.77 25.10 25.27 25.48 25.08 25.81 25.10 24.86 27.17(51%)
VCSUM (zh) 16.36 15.70 15.29 15.03 15.70 15.77 15.59 15.89 15.32 14.67 16.24(28%)
TREC 72.50 72.50 71.00 65.50 60.00 66.50 63.50 70.50 59.50 64.00 73.50(44%)
TriviaQA 91.15 90.41 91.44 90.25 91.47 90.97 91.63 91.14 90.98 81.17 92.14(40%)
SAMSum 43.72 42.68 43.41 42.57 44.21 43.44 43.00 43.41 43.91 39.64 42.61(30%)
LSHT 46.50 46.50 40.00 34.50 45.00 46.00 45.00 46.50 45.00 24.00 45.5(31%)
Passage Count 6.63 6.67 6.73 2.21 6.33 6.07 6.03 4.74 8.04 1.80 9.78(37%)
PassageRetrieval-en 97.98 98.55 96.67 93.54 95.98 97.88 95.83 98.12 96.10 17.96 97.32(36%)
PassageRetrieval-zh 77.99 75.58 43.61 67.83 79.66 78.27 72.38 78.22 79.48 19.72 78.17(34%)
LCC 54.10 56.22 53.10 55.61 54.08 53.13 54.23 52.33 51.87 52.28 53.27(57%)
RepoBench-P 51.39 57.54 50.80 55.17 52.49 51.59 52.32 52.38 51.59 48.46 52.68(47%)
Table 5: LongBench Results with 50% Budget Size on LLama 3.1 8B .
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
NarrativeQA 29.21 26.95 28.76 24.00 24.65 25.40 25.66 28.20 24.41 23.58 28.1(31%)
Qasper 41.34 35.65 35.80 27.88 31.68 31.94 31.38 38.10 31.50 28.60 41.75(46%)
MultiFieldQA-en 52.51 52.50 37.35 40.15 46.12 48.52 44.73 52.20 45.70 35.52 52.79(42%)
MultiFieldQA-zh 58.03 55.37 36.43 39.43 40.55 40.92 43.70 47.93 38.24 27.48 56.75(38%)
HotpotQA 49.62 52.94 47.66 47.24 49.34 49.57 49.51 47.20 51.26 39.64 50.73(39%)
2WikiQA 40.01 39.26 38.44 38.09 37.27 36.75 40.24 38.13 37.49 31.42 40.77(42%)
Musique 28.44 29.45 27.24 20.61 26.43 25.77 27.28 28.54 25.85 16.84 28.64(38%)
DuReader (zh) 34.87 36.44 18.63 26.82 30.29 32.18 32.72 32.18 29.71 20.79 34.7(33%)
GovReport 34.94 32.07 30.79 30.58 31.43 31.55 32.16 32.39 29.63 25.80 34.23(38%)
QMSum 25.67 24.20 22.91 22.48 23.16 24.36 23.47 25.12 24.11 21.08 25.94(37%)
MultiNews 27.85 26.97 25.18 25.62 25.97 26.24 25.83 26.26 25.56 24.39 27.47(54%)
VCSUM (zh) 16.34 15.58 16.24 15.80 15.94 16.42 15.73 16.57 15.63 14.13 16.55(31%)
TREC 75.50 74.50 74.00 66.00 66.50 68.50 69.00 74.50 64.50 69.00 75.00(49%)
TriviaQA 88.89 87.37 89.06 88.36 88.71 88.81 88.69 88.08 89.55 87.05 89.3(41%)
SAMSum 47.29 44.85 47.30 45.82 47.19 47.30 46.77 47.31 46.85 45.29 46.87(35%)
LSHT 39.75 37.50 33.00 31.50 39.25 39.25 39.25 40.25 38.00 21.00 38.75(32%)
Passage Count 5.50 6.00 5.50 4.00 4.00 5.50 5.00 5.00 4.00 4.50 7.50(39%)
PassageRetrieval-en 98.00 99.00 89.50 81.00 97.50 98.50 97.00 98.00 98.50 33.00 98.0(35%)
PassageRetrieval-zh 96.50 96.50 51.50 76.00 89.75 95.50 92.00 97.00 96.50 17.50 96.0(35%)
LCC 53.02 53.22 52.94 54.40 54.50 54.14 53.79 54.63 53.78 52.04 53.66(56%)
RepoBench-P 56.83 55.89 55.47 57.55 56.76 57.05 55.68 56.38 56.36 54.67 57.38(52%)
Table 6: LongBench Results with 50% Budget Size on Mistral-7B. .
Tables  5 and  6 show the evaluation results on LongBench with 50% budgets. Despite having fewer KV cache budgets in all the datasets,  NAtS still achieves comparable performance on the LLama3-8B model and better results on the Mistral model and generally performs comparable to the results with the full attention transformers.
Appendix G Ablation Study
G.1 Sparse Regularization Term  $\lambda$
Full NAtS 1e-6 NAtS 5e-7 NAtS 3e-7 NAtS 1e-7 NAtS 5e-8
NarrativeQA 31.4 27.05(5%) 29.19(9%) 30.53(13%) 31.39(29%) 31.94(41%)
Qasper 24.7 31.39(15%) 28.6(21%) 23.76(27%) 25.17(45%) 26.97(58%)
MultiFieldQA-en 29.5 26.84(13%) 27.41(18%) 28.94(24%) 28.94(41%) 29.7(54%)
MultiFieldQA-zh 60.0 51.73(14%) 56.64(17%) 61.18(22%) 61.27(37%) 61.87(49%)
HotpotQA 17.1 15.47(10%) 17.44(15%) 16.06(20%) 17.28(37%) 16.65(49%)
2WikiQA 16.6 15.37(13%) 16.44(18%) 17.05(23%) 16.84(41%) 16.75(53%)
Musique 11.6 10.78(9%) 9.87(13%) 11.42(18%) 11.89(36%) 11.31(48%)
DuReader (zh) 35.6 30.06(9%) 33.63(13%) 34.28(17%) 33.36(32%) 34.41(45%)
GovReport 34.3 30.82(10%) 32.47(14%) 33.82(19%) 34.23(36%) 34.8(50%)
QMSum 23.3 23.37(8%) 23.09(12%) 23.11(17%) 22.6(33%) 22.74(45%)
MultiNews 27.1 26.45(22%) 26.57(28%) 26.72(33%) 27.17(51%) 27.11(64%)
VCSUM (zh) 16.4 15.56(7%) 15.61(10%) 15.61(14%) 16.24(28%) 15.99(39%)
TREC 72.5 68.0(14%) 70.5(20%) 72.0(26%) 73.5(44%) 73.0(56%)
TriviaQA 91.2 91.38(12%) 90.32(17%) 91.61(22%) 92.14(40%) 91.64(52%)
SAMSum 43.7 43.98(8%) 44.0(12%) 44.21(16%) 42.61(30%) 43.84(41%)
LSHT 46.5 46.0(8%) 44.0(12%) 47.5(16%) 45.5(31%) 45.5(43%)
Passage Count 6.6 2.56(10%) 6.88(14%) 7.12(20%) 9.78(37%) 8.82(50%)
PassageRetrieval-en 98.0 48.17(9%) 82.65(14%) 95.81(19%) 97.32(36%) 96.93(48%)
PassageRetrieval-zh 78.0 33.14(12%) 64.77(16%) 78.5(20%) 78.17(34%) 77.7(46%)
LCC 54.1 55.46(24%) 54.53(31%) 53.54(38%) 53.27(57%) 54.4(69%)
RepoBench-P 51.4 50.78(14%) 51.85(21%) 53.48(28%) 52.68(47%) 52.36(59%)
Table 7: Ablation Study of Sparse Regularization values $\lambda$ for LLama3.1-8B
We first study the impact of sparse regularization terms $\lambda$. This value controls the efficient KV values cached in the model and thus the model performance. The result is shown in Table 7. We underline the results that are better than the optimal baselines with $25\%$ budgets, and bold the results that are better than the optimal baselines with $50\%$. Despite that NAtS in Table 2 (NAtS 3e7NAtS{}\ 3e-7) used more than $25\%$ overall KV budgets for some tasks, here we show NAtSNAtS{} could still outperform many of the corresponding optimal baselines with an even smaller KV budget.
G.2 Sliding Window Length W
Another important hyperparameter for NAtS is the sliding window size W. We apply different sliding window sizes W ($64,128,256,512$) to fine-tune the Llama3-8B model (with $\lambda=3e-7$).
Full NAtS 64 NAtS 128 NAtS 256 NAtS 512
NarrativeQA 31.4 21.61(33%) 30.76(15%) 30.53(13%) 28.96(12%)
Qasper 24.7 17.92(34%) 24.1(28%) 23.76(27%) 28.24(29%)
MultiFieldQA-en 29.5 19.37(34%) 29.02(26%) 28.94(24%) 25.87(25%)
MultiFieldQA-zh 60.0 35.49(35%) 60.37(22%) 61.18(22%) 58.08(26%)
HotpotQA 17.1 13.99(33%) 16.41(23%) 16.06(20%) 17.44(18%)
2WikiQA 16.6 11.52(34%) 16.69(26%) 17.05(23%) 16.48(24%)
Musique 11.6 9.42(33%) 11.62(22%) 11.42(18%) 10.89(16%)
DuReader (zh) 35.6 22.26(34%) 35.75(19%) 34.28(17%) 33.91(16%)
GovReport 34.3 26.63(34%) 33.89(22%) 33.82(19%) 33.05(19%)
QMSum 23.3 20.46(33%) 23.52(19%) 23.11(17%) 22.66(16%)
MultiNews 27.1 23.78(36%) 26.73(33%) 26.72(33%) 26.46(40%)
VCSUM (zh) 16.4 13.8(34%) 16.1(15%) 15.61(14%) 15.9(15%)
TREC 72.5 46.5(34%) 73.0(31%) 72.0(26%) 70.5(25%)
TriviaQA 91.2 84.44(34%) 91.89(26%) 91.61(22%) 91.66(21%)
SAMSum 43.7 38.97(34%) 43.62(18%) 44.21(16%) 44.47(17%)
LSHT 46.5 26.5(34%) 47.0(18%) 47.5(16%) 45.5(15%)
Passage Count 6.6 0.0(34%) 5.75(23%) 7.12(20%) 5.66(18%)
PassageRetrieval-en 98.0 7.13(33%) 96.71(23%) 95.81(19%) 85.78(17%)
PassageRetrieval-zh 78.0 5.15(34%) 81.64(21%) 78.5(20%) 70.66(21%)
LCC 54.1 39.13(36%) 54.92(39%) 53.54(38%) 54.15(44%)
RepoBench-P 51.4 36.87(34%) 51.89(32%) 53.48(28%) 51.34(26%)
Table 8: Ablation Study of Sliding Window Length W for Llama3
The result is shown in Table 8. Overall, all the approaches perform similarly. However, a smaller sliding window size generally results in an overall larger KV cache size. A reduced sliding window size would force the model to apply more Global Tokens and Local Tokens to construct the mid-range correlation since this distance cannot be covered by the sliding window tokens. However, as the number of sliding window size further increases, this compression rate might saturate since the remaining tokens might always require a long-range correlation whose distance is much larger than the sliding window size (e.g., these tokens might require the correlation between two tokens whose distances are larger than 1k or even more).
Appendix H KV size Distributions
(a) KV size of LLama on Narrative QA Dataset
(b) KV size of LLama on Multi-News Dataset
Here, we provide more KV distribution results with different datasets and hyperparameters on the longbench dataset.
Figrues  10(a) and  10(b) show the KV cache size distributions for Llama 3.1 8B on the Narrative QA and Multi News Datasets. The models tend to preserve more KV caches in the shallower layers and ignore the KV caches in the intermediate layers. However, the model still preserves several KV caches near the output layers. Additionally, even in the same layer, the distributions of the KV cache sizes are not evenly distributed in each layer. Some heads are always preferred while the others might be dropped as the context or $\lambda$ changes.
(c) KV size of Mistral on NarrativeQA Dataset
(d) KV size of Mistral on Multi-News Dataset
Figures  10(c) and  10(d) show the KV cache size distributions from the Mistral model. Although the two models share some similar trends, e.g., both models tend to preserve more tokens in the shallower and the last few layers. However, compared to Llama, the KV caches for the Mistral model are more unevenly distributed and tend to gather towards some specific heads. This highlights the importance of adapting different KV eviction strategies to different models.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In this paper, we present NAtS in Section 3. The experimental results in Section 4 show that its strong performance and efficiency in different benchmarks.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes, we discuss the limitations in Section  A.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include theoretical results
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The detailed experiments set up and hyperparameter configurations are described in Section  4 and  E. The search space is described in Section 3 and the algorithm is described in Section D. We also provide the codes under https://github.com/automl/NeuralAttentionSearch
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Yes, all the datasets used in this paper are open-sourced. We further provide a script on how to generate the training set for fine-tuning an LLM with NAtS under: https://github.com/automl/NeuralAttentionSearch
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, we provide the detailed experiments setup and hyperparemter configurations in Section 4 and  E
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: All the experiments in this paper run only once with a fixed random seed.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes, we report the hardware and resources we used to train (4 Nvidia H100 PCIe GPUs for 18 hours) /fine-tune the model (2 Nvidia H100 PCIe GPUs in 8 hours) in Section 4 and  E.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we have reviewed the NeurIPS Code of Ethics.  NAtS is a methodology paper and does not have ethical concerns.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Yes, we have discussed the broader impact in Section B
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes, the data set and models used in our paper are open sourced and we cite the original paper/github that provides the data and model in this paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not release new assets
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The experiments parts involve fine tuning an existing LLM model. We used  NAtS to fine tune a Meta-Llama-3.1-8b-Instruct and Mistral-7B-Instruct-v0.3 model.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.